OpenStudy (anonymous):

A $5,400.00 principal earns 6% interest, compounded annually. After 8 years, what is the balance in the account? Explain how you got your answer, please.

OpenStudy (anonymous):

FV = PV (1+r)^n future value = 5400(1+0.06)^8 FV = $8606.78

OpenStudy (anonymous):

Sorry, I'm not really understanding that @Light&Happiness

OpenStudy (anonymous):

Oh wait, I got it now.